# Multilayer Artificial Benchmark for Community Detection (mABCD)

### v1

@@ 5 Properties of the mABCD Model
Correlations Between Nodes Degrees
@@
To see how quickly the process converges, we compare the results for the number nodes $n$ in $\{10^{3},10^{4},10^{5},10^{6}\}$ and all actors being active—see Figure 4 (Left). To see how close the coefficients are to their asymptotic limit (as $n\to\infty$), we also look at the difference between the values for $n\in\{10^{3},10^{4},10^{5}\}$ and the ones for $n=10^{6}$—see Figure 4 (Right). The differences are rather small, even for $n$ as small as $1{,}000$. Of course, there are some natural fluctuations, but they are rather insignificant compared to the expectation. It justifies our design of the algorithm. In practice, based on the experiments with $n=10^{6}$ (that are saved and accessible by the algorithm), one can select an appropriate value of the parameter $\sigma$ that is expected to produce the desired Kendall rank correlation coefficients $\tau$. With such a choice of $\sigma$, generating 20 independent sequences is enough to make sure that one of them is very close to the desired value of $\tau$.
Figure 4: Left: Kendall rank correlation coefficients between actors’ labels and generated random variables $X_{a}=N(a/n,\sigma)$ for various values of $n$. Right: Difference between Kendall rank correlation coefficients for small values of $n$ and the ones for $n=1{,}000{,}000$.
In the second experiment, we generate the degree sequences for $n=1{,}000$ actors in 5 layers. The desired correlations between the actors’ labels and the order of the corresponding random variables $X_{a}$ is fixed to be $(\tau_{i})=(1.0,0.5,0.0,-0.5,-1.0)$ and not all the actors are active in all layers: $(q_{i})=(1.0,0.9,0.8,0.7,0.6)$. The degree sequence for active nodes is independently generated in each layer with the following parameters: $\gamma_{i}=2.5$, $\delta_{i}=5$, $\Delta_{i}=50$.
We repeat the above experiment 100 times. The experimental means and standard deviations of the corresponding Kendall $\tau$ correlations between the hidden labels of active actors and generated random variables $X_{a}$ are $(1.0,0.5,0.0,-0.5,-1.0)$ and, respectively, $(0.0,0.001,0.002,0.001,0.0)$, very close to the desired sequence $(1.0,0.5,0.0,-0.5,-1.0)$. We also computed correlations between sequences of $X_{a}$ as well as degree sequences in all $\binom{5}{2}$ pairs of layers (on nodes that are active in both layers). The experimental means are reported in Figure 5 for sequences of $X_{a}$ (Left) and for degree sequences (Right). As expected, both matrices contain similar values, and correlations behave as expected. For example, layers 4 and 5 are negatively correlated with the labels of actors ($-0.5$ and $-1.0$, respectively) and so are positively correlated with each other ($0.5$ and $0.537$ in the two corresponding correlations).
@@
The third and final experiment in this subsection is merely a sanity check, as this property is enforced in the model by design. We generate degree sequences for three layers with $n=100{,}000$ actors in each, all of them being active ($q=1$). The minimum and the maximum degrees are fixed to be $\delta=5$, $\Delta=316\approx\sqrt{n}$ across the three layers, but the power-law exponents vary: $(\gamma_{i})=(2.2,2.5,2.8)$. For a given integer $k$, let $f(k)$ be the experimental cumulative degree distribution, that is, $f(k)$ is the fraction of nodes of degree at least $k$. For a given set of parameters, the theoretical cumulative degree distribution is equal to
@@
[…]
Correlations Between Partitions
@@
Figure 8: The average value of AMI (over 100 repetitions) between two partitions generated with $q_{1}$ and, respectively, $q_{2}$: $n=1{,}000$, $s_{1}=s_{2}=8$, $S_{1}=S_{2}=32$, $\beta_{1}=\beta_{2}=1.5$; $d$-dimensional reference layer was used: (left) $d=1$, (right) $d=2$. Only common active nodes are kept in both layers.
Figure 9: The average value of AMI (over 500 repetitions) between two partitions generated with $S_{1}$ and, respectively, $S_{2}$: $n=1{,}000$, $s_{1}=s_{2}=8$, $\beta_{1}=\beta_{2}=1.5$; $d$-dimensional reference layer was used: (left) $d=1$, (right) $d=2$.
In the second experiment, we check how the correlation strength parameter $r$ affects the AMI between the two layers — see Figure 10. It was introduced to provide a smooth transition between the maximum possible correlation (with $r=1$) and completely random partitions (with $r=0$). Our experiments confirm this desired property and show that the AMI is quite stable even for small networks on $n=1{,}000$ nodes.
Figure 10: The average value of AMI $\pm$ 1 standard deviation (over 10 repetitions) between two partitions generated with different values of $r$ ($x$-axis): $n=1{,}000$, $\beta_{1}=\beta_{2}=1.5$; $d$-dimensional reference layer was used: (left) $d=1$, (right) $d=2$.
In the third experiment, we generate partitions of $n=1{,}000$ actors in 5 layers. As before, not all the actors are active in all layers: $(q_{i})=(0.6,0.7,0.8,0.9,1.0)$. We test different values for the distribution of the community sizes: $(\beta_{i})=(1.1,1.3,1.5,1.7,1.9)$, $(s_{i})=(8,24,40,56,72)$, and $(S_{i})=(32,48,64,80,96)$. More importantly, we test different values of the correlation strength between communities and the reference layer: $(r_{i})=(0.0,0.25,0.5,0.75,1.0)$. The experiment is repeated 100 times, and the experimental means of the AMI between $\binom{5}{2}$ pairs of partitions are reported in Figure 11. As expected, the partition of active actors in the first layer (with $r_{1}=0.0$) is not correlated with any other partition. Moreover, the values of AMI increase together with the corresponding correlation strengths.
@@
In the fourth and final experiment in this subsection, we generate sequences of community sizes with different distributions. There are $n=100{,}000$ actors in each of the three layers, all of them being active ($q=1$). The minimum and the maximum community size is fixed to be $s=10$ and $S=1000$ across the three layers, but the power-law exponents vary: $(\beta_{i})=(1.2,1.5,1.8)$. We show that the experimental community sizes distributions are very close to the desired ones — see Figure 6 (Right) for the cumulative community sizes distributions of the three layers.
Correlations Between Edges in Various Layers
@@
Speed
@@
One of the reasons why the original ABCD model was developed was the fact (pointed out by the industry partners we worked with back then) that other synthetic random graph models with community structure are inherently slow. Keeping this concern in mind, the design of the ABCD model aims to generate networks fast. Indeed, after careful preprocessing (that is very fast), ABCD is essentially a union of independent copies of the configuration model. In two minutes, on a typical laptop, one can generate networks on $n=1.5\cdot 2^{21}$ nodes. The experiments reported in the section were executed on Macbook Air M1 (2020) with macOS v15.5.
@@
The mABCD model is clearly much more complicated than the original ABCD, but it builds on fast ingredients of ABCD. Not surprisingly, Phase 6 (switching edges to get the desired correlations between edges in various layers) is the most time-consuming phase. Moreover, it is clear that the more layers, the longer the process is. In two minutes, one can generate mABCD networks on $n=2^{16}=65{,}536$ nodes. However, without Phase 6, in the same time one can produce much larger networks on $n=2^{20}=1{,}048{,}576$ nodes — see Figure 13. The percentage breakdown is presented in Figure 14.
@@
Finally, we have compared the speed of mABCD model to the multilayerGM framework [6]; see Section 2.1 for more details. The results are presented in Figure 15 and Table 4, and indicate that for the same setup, mABCD is roughly 10 times faster for small networks (1,024 nodes), and the difference in speed quickly increases with the network size (for a two-layer network with 32,768 nodes mABCD is 310 times faster). Since generating one five-layer network with 32,768 nodes took more than one hour for multilayerGM, and we repeated experiments ten times for each combination of node and layer, we stopped our comparison there.
@@ 6.1 Scope and Methodology
Each experiment involves manipulating a single parameter of the mABCD model, with the aim of evaluating networks that differ from the baseline by only that feature. Similarly to the baseline graph, we generated 20 multilayer networks per each set of configuration parameters to mitigate the randomness inherent in mABCD. Diffusion is then simulated on these networks using the same spreading conditions as for the “digital twin” of the real-world network. Since the employed spreading model is also nondeterministic, each simulation was repeated 30 times to reduce the influence of random effects. Finally, the diffusion effectiveness observed in the modified and original networks is compared to assess the influence of the parameter under investigation on the spreading process.
@@
As the influence spreading mechanism, we utilize the Independent Cascade Model [26] (ICM), extended to the multilayer scenario (MICM). This adaptation draws upon a similar rationale to that used in extending the related Linear Threshold Model (LTM), as described by [58]. Consistent with their approach, we regard actors as the principal entities in the diffusion process, while their corresponding nodes across various layers serve a supplementary function. In the original ICM, the probability that an active agent will activate its inactive neighbour is controlled by a single parameter, $\dot{\pi}$, with each activation attempt occurring only once. Our modification introduces an additional parameter, $\dot{\delta}$, termed the protocol function, which dictates how activation signals from different layers are combined to establish the actor’s overall state. Activation of the actor (and all its layer-specific nodes) occurs only if the protocol condition is fulfilled; if not, the actor remains inactive even if some layer-specific nodes receive positive activation signals. We examine two extreme cases for $\dot{\delta}$: the $AND$ protocol, where activation requires positive signals from all layers representing the actor; and the $OR$ protocol, where activation in any single layer suffices.
@@
At this point it is worth noting that, in contrast to the ICM, employing its well-known counterpart in the experimental part, the LTM [28], can lead to unstable results. This is due to the way mABCD generates graphs, namely, it aims to capture community structure (i.e., global properties) rather than local ones. As a result, graphs generated with mABCD contain groups of nodes that induce denser subgraphs compared to the global density, but locally neighbourhoods still resemble tree structure, as there are very few short cycles. By contrast, real networks do not exhibit this local behaviour; they contain many short cycles, even triangles. Consequently, diffusion in mABCD-generated networks under the LTM, which relies on a threshold mechanism requiring a certain number of active neighbours to activate a given node, can either quickly die out or abruptly saturate the network, with no parameter range that produces gradual, stable spreading. On the other hand, the mechanism inherent to the ICM is “immune” to this tree-like local structure, as it focuses on individual edges rather than the entire neighbourhood of a given node.
@@
Finally, the metric used to compare the effectiveness of spreading was gain [18]. It quantifies the total network coverage achieved by initiating the diffusion from a seed set $S_{0}$. Specifically, it is computed as the proportion of actors activated during the diffusion process, excluding the initial seeds, relative to the total number of activatable actors (i.e., those not in $S_{0}$):
@@
The simulations described in this chapter were conducted on a workstation running Ubuntu 20.04.4 LTS with kernel version 6.5.3-arch1-1, equipped with 376 GB of RAM and an Intel(R) Xeon(R) Gold 6238 CPU @ 2.10 GHz (x86_64 architecture). The source code was written in Python 3.12 utilizing mABCD v1.0, networkx v3.3 [29], and network-diffusion v0.18 [19], with particular attention to reproducibility. All scripts, data, and results are publicly available at Github repository [^7] https://github.com/anty-filidor/spreading-vs-mln-structure.

### v2

@@ 5 Properties of the mABCD Model
5.1 Correlations between node degrees
@@
To see how quickly the process converges, we compare the results for the number nodes $n$ in $\{10^{3},10^{4},10^{5},10^{6}\}$ and all actors being active—see Figure 5 (Left). To see how close the coefficients are to their asymptotic limit (as $n\to\infty$), we also look at the difference between the values for $n\in\{10^{3},10^{4},10^{5}\}$ and the ones for $n=10^{6}$—see Figure 5 (Right). The differences are rather small, even for $n$ as small as 1,0001{,}000. Of course, there are some natural fluctuations, but they are rather insignificant compared to the expectation. It justifies our design of the algorithm. In practice, based on the experiments with $n=10^{6}$ (that are saved and accessible by the algorithm), one can select an appropriate value of the parameter $\sigma$ that is expected to produce the desired Kendall rank correlation coefficients $\tau$. With such a choice of $\sigma$, generating 20 independent sequences is enough to make sure that one of them is very close to the desired value of $\tau$.
Figure 5: Left: Kendall rank correlation coefficients between actors’ labels and generated random variables $X_{a}=N(a/n,\sigma)$ for various values of $n$. Right: Difference between Kendall rank correlation coefficients for small values of $n$ and the ones for n=1,000,000n=1{,}000{,}000.
In the second experiment, we generate the degree sequences for n=1,000n=1{,}000 actors in 5 layers. The desired correlations between the actors’ labels and the order of the corresponding random variables $X_{a}$ is fixed to be $(\tau_{i})=(1.0,0.5,0.0,-0.5,-1.0)$ and not all the actors are active in all layers: $(q_{i})=(1.0,0.9,0.8,0.7,0.6)$. The degree sequence for active nodes is independently generated in each layer with the following parameters: $\gamma_{i}=2.5$, $\delta_{i}=5$, $\Delta_{i}=50$.
We repeat the above experiment 100 times. The experimental means and standard deviations of the corresponding Kendall $\tau$ correlations between the hidden labels of active actors and generated random variables $X_{a}$ are (1.0,0.5,0.0,0.5,1.0)(1.0,0.5,0.0,-0.5,-1.0) and, respectively, $(0.0,0.001,0.002,0.001,0.0)$, very close to the desired sequence (1.0,0.5,0.0,0.5,1.0)(1.0,0.5,0.0,-0.5,-1.0). We also computed correlations between sequences of $X_{a}$ as well as degree sequences in all $\binom{5}{2}$ pairs of layers (on nodes that are active in both layers). The experimental means are reported in Figure 6 for sequences of $X_{a}$ (Left) and for degree sequences (Right). As expected, both matrices contain similar values, and correlations behave as expected. For example, layers 4 and 5 are negatively correlated with the labels of actors (0.5-0.5 and 1.0-1.0, respectively) and so are positively correlated with each other ($0.5$ and $0.537$ in the two corresponding correlations).
@@
The third and final experiment in this subsection is merely a sanity check, as this property is enforced in the model by design. We generate degree sequences for three layers with n=100,000n=100{,}000 actors in each, all of them being active ($q=1$). The minimum and the maximum degrees are fixed to be $\delta=5$, $\Delta=316\approx\sqrt{n}$ across the three layers, but the power-law exponents vary: $(\gamma_{i})=(2.2,2.5,2.8)$. For a given integer $k$, let $f(k)$ be the experimental cumulative degree distribution, that is, $f(k)$ is the fraction of nodes of degree at least $k$. For a given set of parameters, the theoretical cumulative degree distribution is equal to
@@
Figure 7: Left: Empirical (dots) and theoretical (dashes) log-log cumulative degree distributions for three degree sequences: $n=100\,000,\delta=5,\Delta=316.$ Right: Empirical (dots) and theoretical (dashes) log-log cumulative community sizes distributions for three sizes sequences: n=100 000,s=10,S=1 000.n=100\,000,s=10,S=1\,000.
5.2 Correlations between partitions
@@
Figure 9: The average value of AMI (over 100 repetitions) between two partitions generated with $q_{1}$ and, respectively, $q_{2}$: n=1,000n=1{,}000, $s_{1}=s_{2}=8$, $S_{1}=S_{2}=32$, $\beta_{1}=\beta_{2}=1.5$; $d$-dimensional reference layer was used: (left) $d=1$, (right) $d=2$. Only common active nodes are kept in both layers.
Figure 10: The average value of AMI (over 500 repetitions) between two partitions generated with $S_{1}$ and, respectively, $S_{2}$: n=1,000n=1{,}000, $s_{1}=s_{2}=8$, $\beta_{1}=\beta_{2}=1.5$; $d$-dimensional reference layer was used: (left) $d=1$, (right) $d=2$.
In the second experiment, we check how the correlation strength parameter $r$ affects the AMI between the two layers — see Figure 11. It was introduced to provide a smooth transition between the maximum possible correlation (with $r=1$) and completely random partitions (with $r=0$). Our experiments confirm this desired property and show that the AMI is quite stable even for small networks on n=1,000n=1{,}000 nodes.
Figure 11: The average value of AMI $\pm$ 1 standard deviation (over 10 repetitions) between two partitions generated with different values of $r$ ($x$-axis): n=1,000n=1{,}000, $\beta_{1}=\beta_{2}=1.5$; $d$-dimensional reference layer was used: (left) $d=1$, (right) $d=2$.
In the third experiment, we generate partitions of n=1,000n=1{,}000 actors in 5 layers. As before, not all the actors are active in all layers: $(q_{i})=(0.6,0.7,0.8,0.9,1.0)$. We test different values for the distribution of the community sizes: $(\beta_{i})=(1.1,1.3,1.5,1.7,1.9)$, $(s_{i})=(8,24,40,56,72)$, and $(S_{i})=(32,48,64,80,96)$. More importantly, we test different values of the correlation strength between communities and the reference layer: $(r_{i})=(0.0,0.25,0.5,0.75,1.0)$. The experiment is repeated 100 times, and the experimental means of the AMI between $\binom{5}{2}$ pairs of partitions are reported in Figure 12. As expected, the partition of active actors in the first layer (with $r_{1}=0.0$) is not correlated with any other partition. Moreover, the values of AMI increase together with the corresponding correlation strengths.
@@
In the fourth and final experiment in this subsection, we generate sequences of community sizes with different distributions. There are n=100,000n=100{,}000 actors in each of the three layers, all of them being active ($q=1$). The minimum and the maximum community size is fixed to be $s=10$ and $S=1000$ across the three layers, but the power-law exponents vary: $(\beta_{i})=(1.2,1.5,1.8)$. We show that the experimental community sizes distributions are very close to the desired ones — see Figure 7 (Right) for the cumulative community sizes distributions of the three layers.
5.3 Correlations between edges in various layers
@@
5.4 Community detection in mABCD networks
As noted in Section 1, the mABCD model supports the validation and assessment of community detection algorithms. Its generation of ground-truth communities enables direct comparison with partitions produced by tested methods. As an extension of ABCD, mABCD incorporates findings from prior analyzes of monoplex community detection algorithms [kaminski2021artificial, kaminski2022modularity]. An open question remains regarding its utility for validating multiplex partitioning algorithms [magnani2021community]. Although relevant, this issue lies beyond the scope of this study. We therefore conduct a simple validation experiment using one multiplex partitioning algorithm and the monoplex counterpart, leaving broader analysis for future work.
We generated a network with 1,000 actors and three layers. Configuration of layers was identical ($q=1,\tau=0.0,\gamma=2.5,\delta=2,\Delta=25,\beta=1.5,s=16,S=64$) except for the community correlation coefficient $r$ ($r_{1}=1.0,r_{2}=0.5,r_{3}=0.0$). The edges correlation matrix $\mathbf{R}$ was not defined, hence no edges rewiring occurred. The noise ratio $\xi$ varied from 0.1 to 0.9 to simulate increasing inter-community connectivity. Due to the stochastic nature of mABCD, ten graphs were produced for each $\xi$. Community detection was performed using the Generalized Louvain algorithm [Mucha_2010] on the whole multiplex network and the base Louvain algorithm [blondel2008fast] on each layer independently. The resulting partitions were compared with the ground truth from mABCD via the AMI metric. The experiment aimed to examine how the quality of group detection depends on two parameters: the fraction of inter-community edges within layers ($\xi$) and the community alignment across layers ($r$). The Generalized Louvain algorithm is capable of detecting pillar communities [magnani2021community]. One can think about pillar communities as groups comprising the same nodes that consistently emerge across the network layers. Therefore, mABCD communities in layers with high $r$ (correlated with latent layer, which implies that also between themselves) should create higher-level structures possible to detect with multiplex algorithm. Generalized Louvain highlights pillar communities by producing consistent clusters’ labels for nodes in all layers. For reporting purposes, we compare the clustering for each layer separately (Figure 14).
Figure 14: The average ($\pm$ 1 std) Adjusted Mutual Information between the ground truth and a) Generalized Louvain’s communities (left), b) Louvain’s communities (right); 50 repetitions for each $\xi$. (Note that for Louvain’s communities all three lines overlap; for Generalized Louvain lines for $r_{1}=1.0$ and $r_{2}=0.5$ are almost overlapped.)
Aligned with expectations, the results show a decline in clustering quality with increasing noise. Clustering within each layer converges to random partitioning for high $\xi$ values. Notably, Generalized Louvain exhibits curious tendency related to the $r$ parameter. The method performed equally well for perfectly correlated ($r_{1}=1.0$) and moderately correlated ($r_{2}=0.5$) layers (the reason is that the algorithm does not see the latent layer, so both layers are indistinguishable from the perspective of community detection, since the third layer is uncorrelated). Performance for the uncorrelated layer is significantly lower, which is expected as the assignment of actors to communities is randomly shuffled, so the signal coming from it to the Generalized Louvain algorithm is weaker. Captured similarities may indicate the ability of Generalized Louvain to detect intra-layer patterns that affect the structure of communities. In contrast, the performance of the Louvain algorithm is much better than the multiplex extension. The monoplex variant is insensitive to the value of $r$, which is aligned with mABCD design. The $r$ parameter controls correlation between layers, but does not affect strength of communities within layers — this property is controlled by the $\xi$ parameter.
5.5 Computational efficiency of the model
@@
One of the reasons why the original ABCD model was developed was the fact (pointed out by the industry partners we worked with back then) that other synthetic random graph models with community structure do not perform well in terms of computational efficiency. Keeping this concern in mind, the design of the ABCD model aims to generate networks fast. Indeed, after careful preprocessing (that is very fast), ABCD is essentially a union of independent copies of the configuration model. In two minutes, on a typical laptop, one can generate networks on $n=1.5\cdot 2^{21}$ nodes. The mABCD model is clearly much more complicated than the original ABCD, but it builds on fast ingredients of ABCD. Therefore, in order to validate whether it preserves computational advantages inherited from its monoplex counterpart, we performed experiments aimed to address this problem. The data reported below was obtained during simulations executed on Macbook Air M1 (2020) with macOS v15.5.
@@
Not surprisingly, Phase 6 (switching edges to get the desired correlations between edges in various layers) is the most time-consuming phase. Moreover, it is clear that the more layers, the longer the process is. In two minutes, one can generate mABCD networks on $n=2^{16}=65{,}536$ nodes. However, without Phase 6, in the same time one can produce much larger networks on $n=2^{20}=1{,}048{,}576$ nodes — see Figure 15. The percentage breakdown is presented in Figure 16.
@@
Finally, we have compared the speed of mABCD model to the multilayerGM framework proposed by [bazzi2020framework] (see Section 2.1 for more details). The results are presented in Figure 17 and Table 4, and indicate that for the same setup, mABCD is roughly 10 times faster for small networks (1,024 nodes), and the difference in speed quickly increases with the network size (for a two-layer network with 32,768 nodes mABCD is 310 times faster). Since generating one five-layer network with 32,768 nodes took more than one hour for multilayerGM, and we repeated experiments ten times for each combination of node and layer, we stopped our comparison there.
@@ 6.1 Scope and methodology
Each experiment involves manipulating a single parameter of the mABCD model, with the aim of evaluating networks that differ from the baseline by only that feature. Similarly to the baseline graph, we generated 20 multilayer networks for each set of configuration parameters to mitigate the randomness inherent in mABCD. Diffusion is then simulated on these networks using the same spreading conditions as for the “digital twin” of the real-world network. Since the employed spreading model is also nondeterministic, each simulation was repeated 30 times to reduce the influence of random effects. Finally, the diffusion effectiveness observed in the modified and original networks is compared to assess the influence of the parameter under investigation on the spreading process.
@@
As the influence spreading mechanism, we utilize the Independent Cascade Model [goldenberg2001icm] (ICM), extended to the multilayer scenario (MICM). This adaptation draws upon a similar rationale to that used in extending the related Linear Threshold Model (LTM), as described by [zhong2022mltm]. Consistent with their approach, we regard actors as the principal entities in the diffusion process, while their corresponding nodes across various layers serve a supplementary function. In the original ICM, the probability that an active agent will activate its inactive neighbour is controlled by a single parameter, $\dot{\pi}$, with each activation attempt occurring only once. Our modification introduces an additional parameter, $\dot{\delta}$, termed the protocol function, which dictates how activation signals from different layers are combined to establish the actor’s overall state. Activation of the actor (and all its layer-specific nodes) occurs only if the protocol condition is fulfilled; if not, the actor remains inactive even if some layer-specific nodes receive positive activation signals. We examine two extreme cases for $\dot{\delta}$: the $AND$ protocol, where activation requires positive signals from all layers representing the actor; and the $OR$ protocol, where activation in any single layer suffices.
@@
At this point it is worth noting that, in contrast to the ICM, employing its well-known counterpart in the experimental part, the LTM [granovetter1978threshold], can lead to unstable results. This is due to the way mABCD generates graphs, namely, it aims to capture community structure (i.e., global properties) rather than local ones. As a result, graphs generated with mABCD contain groups of nodes that induce denser subgraphs compared to the global density, but locally neighbourhoods still resemble tree structure, as there are very few short cycles. By contrast, real networks do not exhibit this local behaviour; they contain many short cycles, even triangles. Consequently, diffusion in mABCD-generated networks under the LTM, which relies on a threshold mechanism requiring a certain number of active neighbours to activate a given node, can either quickly die out or abruptly saturate the network, with no parameter range that produces gradual, stable spreading. On the other hand, the mechanism inherent to the ICM is “immune” to this tree-like local structure, as it focuses on individual edges rather than the entire neighbourhood of a given node.
@@
Finally, the metric used to compare the effectiveness of spreading was gain [czuba2024rankrefininginfmaxmln]. It quantifies the total network coverage achieved by initiating the diffusion from a seed set $S_{0}$. Specifically, it is computed as the proportion of actors activated during the diffusion process, excluding the initial seeds, relative to the total number of activatable actors (i.e., those not in $S_{0}$):
@@
The simulations described in this chapter were conducted on a workstation running Ubuntu 20.04.4 LTS with kernel version 6.5.3-arch1-1, equipped with 376 GB of RAM and an Intel(R) Xeon(R) Gold 6238 CPU @ 2.10 GHz (x86_64 architecture). The source code was written in Python 3.12 utilizing mABCD v1.0, networkx v3.3 [hagberg2008networkx], and network-diffusion v0.18 [czuba2024networkdiffusion], with particular attention to reproducibility. All scripts, data, and results are publicly available at Github repository [^8] https://github.com/anty-filidor/spreading-vs-mln-structure.